OpenStudy (anonymous):

solve the following system of equations x +5y = 7 (1) x=3-5y (2) i dont know how to do this becasue for ()s

OpenStudy (anonymous):

that means equation 1 and equation 2 they are not in the real equation.

OpenStudy (anonymous):

there are no solutions the lines are parallel

OpenStudy (anonymous):

y=-1/5 x+7 and y=-1/5x+3 will not cross, therefore no solution exists, you can also see this if you subtract the equations, 0 = 4 doesn't work